OpenStudy (anonymous):

Julianne wants to take a trip around the world. She plans to deposit $175 at the beginning of each month into an investment with a 3.25% interest rate, compounded monthly. How much will she have in the account after 12 years? $30,770.00 $30,853.33 $533,286.11 $550,617.91

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

this question has to do with annuities. You're making regular deposits into a savings account every month for 12 years. Your money earns interest. We want to know how much money you'll have in this account after that length of time. do you have available a reference that presents various formulas for annuities? I have probably done this kind of problem before, but not for many years, so I don't remember the correct formula offhand. If you have this info in online materials, perhaps you could share the URL with me, so that we could decide which formula to use.

ganeshie8 (ganeshie8):

lookup for "annuity due" formula in ur notes

OpenStudy (mathmale):

See http://www.investopedia.com/articles/03/101503.asp This web page is a bit too advanced, but I'm sure the proper formula could be found there. I did a search for "annuity formulas." You could do your own search for "annuity formulas," or do as ganeshie8 suggests and look for anything connected to "annuities" in your notes or online learning materials.

OpenStudy (mathmale):

Note that you are looking for the FUTURE VALUE of a long string of payments INTO an annuity. What will all these funds amount to after 12 years of paying in something every month?

ganeshie8 (ganeshie8):

FV of Annuity Due = \(\large \left[ C\bullet \dfrac{(1+i)^{nt}-1}{i}\right](1+i)\)

ganeshie8 (ganeshie8):

Does this formula looks familiar ?

ganeshie8 (ganeshie8):

just figure out the variables and plugin : \(C\) is periodic cash payment you must have seen other variables before :)

OpenStudy (anonymous):

Yes it does that's what i was looking for in my notes except mine doesnt say FV mine says \[PVOA= C (\frac{ 1 }{ i } - \frac{ 1 }{ i(1+i)^{nt-1} }\]

ganeshie8 (ganeshie8):

dont use that, thats a different formula

ganeshie8 (ganeshie8):

look for \(FVAD\) Future Value of Annuity Due

ganeshie8 (ganeshie8):

\(\large FVAD = \left[ C\bullet \dfrac{(1+i)^{nt}-1}{i}\right](1+i)\)

ganeshie8 (ganeshie8):

^should look like that...

OpenStudy (anonymous):

I have PVAD that looks like the first one i typed but its adding C at the end i dont have anything with future value of annuity

ganeshie8 (ganeshie8):

PVAD is not useful here

ganeshie8 (ganeshie8):

you must use FVAD formula

OpenStudy (anonymous):

how is it 30853? @ganeshie8

ganeshie8 (ganeshie8):

plug the numbers in FVAD formula, and evaluate

OpenStudy (anonymous):

i did and i got 550,617

OpenStudy (anonymous):

oh nevermind i didnt divide .0325 by 12 @ganeshie8

ganeshie8 (ganeshie8):

lol i do that mistake everytime !

ganeshie8 (ganeshie8):

basically you need to remember these 4 formulas when working with annuities : 1) FVOD 2) FVAD 3) PVOD 4) PVAD

OpenStudy (mathmale):

@ganeshie8: Great work! Care to share any online references you may know of in regard to 1) FVOD 2) FVAD 3) PVOD 4) PVAD ??

ganeshie8 (ganeshie8):

http://web.calstatela.edu/faculty/dhossai2/common/resources/TABLES%20PV_FV/ first four tables list these formulas; deriving these is bit easy.... i had worked these and other monthly payment formulas couple of weeks back... took more than a week to fully understand them though... they just give geometric series... pretty neat stuff :)

ganeshie8 (ganeshie8):

I dont use PV* formulas actually... I only use FV* formulas and scale the amount for finding PV...

OpenStudy (mathmale):

I surely have to admire people like yourself who can come up with their own formulas instead of relying on a table or other reference. :)